# “Sharing, Not Showing Off”: How BeReal Approaches Authentic Self-Presentation on Social Media Through Its Design

### v1

@@ Abstract.
Adolescents are particularly vulnerable to the pressures created by social media, such as heightened self-consciousness and the need for extensive self-presentation. In this study, we investigate how BeReal, a social media platform designed to counter some of these pressures, influences adolescents’ self-presentation behaviors. We interviewed 29 users aged 13-18 to understand their experiences with BeReal. We found that BeReal’s design focuses on spontaneous sharing, including randomly timed daily notifications and reciprocal posting, discourages staged posts, encourages careful curation of the audience, and reduces pressure on self-presentation. The space created by BeReal offers benefits such as validating an unfiltered life and reframing social comparison, but its approach to self-presentation is sometimes perceived as limited or unappealing and, at times, even toxic. Drawing on this empirical data, we distill a set of design guidelines for creating platforms that support authentic self-presentation online, such as scaffolding reciprocity and expanding beyond spontaneous photo-sharing to allow users to more accurately and comfortably portray themselves.
@@ 1. Introduction
The number of people who report having no one to discuss difficult subjects with has tripled in recent decades, and the average person’s social network has shrunk by one-third (McPherson et al., 2006). A significant increase in social isolation has occurred, with nearly 25% of Americans in 2004 reporting they had no confidant, up from 10% in 1985 (McPherson et al., 2006). This is considering the profound importance of social relationships on our physical and mental health, longevity, and happiness (Cacioppo and Cacioppo, 2014; House et al., 1988; Cohen, 2004). The Study of Adult Development examined participants over 75 years and indicates that good relationships keep us happier and healthier (Waldinger and Schulz, 2016). A 2010 review of over 300,000 participants found that weak social ties are as detrimental to health as alcoholism (Holt-Lunstad et al., 2010). For adolescents, this intrinsic need for trusted interpersonal connection, particularly with peers, is even more pronounced than with adults (Erikson, 1994a; Davis, 2012). And with the widespread adoption of social media, the nature of these social interactions has transformed: 95% of adolescents ages 13-17 engage actively with friends through social media platforms (noa, 2023). As these peer relationships increasingly have an online dimension, it is crucial to assess how technology supports teens’ relationship-building.
@@
Social Penetration Theory (Carpenter and Greene, 2015a) explains that relationships develop through reciprocal and authentic self-disclosure, which progresses in depth over time. However, social media often offers “empty calories of social interaction” (Prinstein, 2023), which might seem to fulfill adolescents’ developmental needs, but actually lack the substantive qualities necessary for true social benefits. The use of these platforms can amplify inherent developmental vulnerabilities of adolescents, such as increased sensitivity to social feedback and emotional reactivity due to ongoing brain maturation (Somerville, 2013a; Steinberg, 2005; Falk et al., 2014; Peake et al., 2013; Somerville, 2013b), which can further detract them from forming meaningful and supportive relationships. Nonetheless, a majority of teens use social media and regularly engage in extensive self-presentation through manicured content and the anxiety and distorted perception of reality that it can provoke (Aacap, [n. d.]).
@@
BeReal is a social media platform created to combat the difficulty of cultivating relationships online due to the sparsity of authentic interactions. BeReal has a specific focus: “everyday get a notification to share a genuine glimpse into your real life with the people you care most about” which they believe enables users to “show your friends who you really are” (noa, [n. d.]). Though it was launched in 2019, long before the surgeon general’s advisory, its design and mission align with the surgeon general’s recommendation that tech companies “seek to maximize the potential benefits and avoid design features that attempt to maximize time, attention, and engagement” (noa, 2023).
@@
To answer these questions, we conducted semi-structured interviews with 29 BeReal users aged 13 to 18, asking about their general experiences with BeReal. We found that BeReal’s features and affordances are effective in encouraging users to share casual, authentic snapshots of their daily lives. Features such as randomly timed daily notifications encourage authentic self-presentation via unstaged, unfiltered photo-sharing, and design elements like reciprocal posting encourage users to curate their audience, thereby making their self-presentation process less intentional or controlled. The participants saw many benefits of such a design. They especially appreciated that BeReal nudged users to collectively share more casual and less curated content, a change they said they sought on social media. However, some participants also felt that their ability to present themselves accurately and intentionally was too limited, and they explained that unfiltered photos taken at unplanned times are just one way of giving others insight into their lives. Others said that BeReal’s explicit prioritization of “authenticity” in self-presentation sometimes led to toxic competition and judgment about which users are truly being authentic.
In this paper, we provide empirical evidence of BeReal’s success in encouraging adolescent users to present themselves authentically. This serves as a proof of concept that design can significantly influence self-presentation. We also document the limitations of BeReal’s approach and outline design recommendations for creating a platform that encourages developmentally supportive online self-presentation for adolescents, such as scaffolding reciprocal sharing and supporting self-presentations that actors find to be accurate portrayals of themselves. Based on our findings, we introduce the Social Media Self-Presentation Matrix (SMSM). Grounded in theories on self-presentation goals and processes (Schlenker, 2003), this matrix presents ways that social media design can support different self-presentation strategies. It does so through four facets of self-presentation, spanning axes from automatic to controlled and from self-enhancing to self-expressive. Using the matrix, we highlight an untapped design opportunity to support controlled and self-expressive self-presentation, which could offer an alternative way for social media platforms to achieve the goals initially intended by BeReal. Further, we propose a design agenda to support better relationship-building online based on authenticity for teens on social media. We hope this work will encourage designers and researchers to more comprehensively support authentic exchanges in teen peer-relationship-building on social media, addressing a developmental need that is currently underrepresented.
@@ 2.1. Adolescents and the Developmental Significance of Authentic Self-Disclosure
Forming close, trusting relationships is fundamental to people’s sense of meaning, happiness, and long-term well-being (Argyle, 2013). For adolescents, this intrinsic need for trusted interpersonal connections, particularly with peers, is even more pronounced than it is for adults (Erikson, 1994b; Davis, 2012). As described by the Social Penetration Theory (SPT) (Carpenter and Greene, 2015b), interpersonal relationships develop from initial, surface-level interactions to deeper, more intimate connections through a gradual process of self-disclosure. The breadth and depth of these disclosures increase as the relationship strengthens and matures. Also central to this progression is the norm of reciprocity, which posits that mutual exchange is fundamental to these evolving relationships.
@@
For self-disclosure to effectively enhance intimacy and trust in relationships, it must be authentic. Authenticity—the characteristic of being in accordance with one’s internal values than with external contingencies (Peets and Hodges, 2018)—ensures that the information shared truly represents the individual’s self, fostering genuine connections and understanding between people (Laurenceau et al., 1998; Peets and Hodges, 2018). Moreover, expressing their authentic selves and receiving acknowledgment and affirmation of these selves from peers is crucial for individuals, particularly adolescents, to achieve a clear and coherent self-concept (Gardner, 2013). The formation of a coherent identity, rooted in an authentic self that is recognized and validated by one’s social circle, is fundamental to adolescent development (Erikson, 1994a). Research has also shown that authenticity in self-disclosure is associated with life satisfaction (Bailey et al., 2020), particularly among girls, linking authenticity to self-esteem and overall well-being (Impett et al., 2008). Moreover, a higher degree of authenticity in self-disclosing communication is connected to enhanced self-perceptions, reduced feelings of isolation in school, and greater satisfaction in relationships among adolescents (Peets and Hodges, 2018).
@@
In recent years, particularly exacerbated by the COVID-19 pandemic and subsequent physical distancing measures, adolescents have increasingly turned to social media for social connection (Hamilton et al., 2022; Pitt et al., 2021). As of 2022, nearly all teenagers aged 13-17 (95%) own a smartphone or access the internet (97%), with nearly half (46%) using the internet ‘almost constantly’ (Vogels et al., 2022). Previous studies have identified that the primary motivation behind teens’ use of digital media is to maintain connections with existing friends (Ito, 2013; Boyd, 2008a; Davis, 2012). Given the critical role of online platforms in adolescent lives and the value placed on authenticity in their relationships, it is understandable that modern teenagers actively seek exchanges of authentic self-disclosure on social media and broader media contexts (Darr and Doss, 2022; Stephanie Rivas-Lara Becky Pham Jasmine Baten Annie Meyers Yalda T. Uhls, 2022). Further, the significantly reduced social cues on online platforms heighten uncertainty, making self-disclosure especially crucial for (adolescent) users to understand each other and strengthen their connections (Schlosser, 2020; Tidwell and Walther, 2002). This underscores the importance of examining how technology can support or undermine the development and maintenance of these relationships, with a particular focus on the reciprocal exchange of authentic self-disclosure throughout the stages of social penetration.
@@ 2.2. Authentic Self-Disclosure and Self-Presentation
Given that self-disclosure, especially in online contexts, involves extensive interpersonal contexts is social behavior that is likened a performance (Goffman, 2016) that symbolically communicates information about self to others (Blumer, 1986), self-presentation readily happens. Self-disclosure is defined as “verbally communicating personal information about the self to another person” (Forgas, 2011) (p. 449), and reflects communicating a factual representation of oneself, regardless of its effect on one’s public self-image. In contrast, self-presentation is defined as ‘‘the goal-directed activity of controlling information to influence the impressions formed by an audience about the self” (Schlenker and Wowra, 2003) (p. 871). Thus, self-disclosure typically aligns with authenticity, whereas self-presentation may suggest a lack of authenticity. However, these two are not mutually exclusive (Johnson, 1981); revealing true aspects of oneself does not necessarily equate to self-disclosure, as such revelations can also support a favorable public persona (Schlosser, 2020). Conversely, self-presentation should not be dismissed as “just superficial, deceptive, or manipulative” (Schlenker, 2003). In summary, we adopt the view in this paper that self-presentation is a strategic form of self-disclosure that predominates on social media, where users, acting as performers, anticipate positive outcomes from their disclosures. However, this form of self-presentation often conflicts with authenticity unless the self-presentation itself is authentic.
@@
Although self-presentation is often perceived as strategic gamesmanship, this view does not capture its full complexity (Schlenker, 2003). In reality, self-presentation can be deeply “authentic,” as in, it typically involves a portrayal that is somewhat enhanced and idealized, yet genuinely believed by the person presenting it (Brown, 2014; Greenwald and Breckler, 1985; Williams and Gilovich, 2008). Sometimes the goal of self-presentation is accurate self-portrayal (Baumeister, 1982; Cheek and Hogan, 1983; Leary, 2019; Schlenker, 1980; Schlenker and Pontari, 2000). The aim could be to ensure that others perceive one in a manner that secures the respect and treatment aligned with one’s identity, to receive affirming feedback that addresses personal uncertainties, or to embody the principle that “honesty is the best policy,” thereby feeling authentic while reducing the risks associated with deception. Furthermore, intentional self-packaging does not necessarily indicate inauthenticity; akin to how a skilled lecturer carefully crafts their presentation to effectively convey genuine and significant information to their audience—contrasting starkly with a disorganized lecturer whose lack of preparation may seem insincere and self-absorbed—the careful arrangement of personal information can reflect an authentic depiction of oneself (as seen by the individual), not merely a false facade (Schlenker and Pontari, 2000). In other words, presenting accurate information requires as much skill in self-presentation as presenting inaccurate information if it is to achieve the intended effect on the audience.
@@
Authentic self-presentation can manifest through either automatic or controlled cognitive processes. Automatic processes are characterized by 1) occurring without conscious awareness, 2) requiring minimal cognitive effort, 3) being autonomous and not requiring conscious monitoring, and 4) being initiated involuntarily without outside triggers (Bargh, 1996). In such instances, self-presentation scripts may be triggered automatically by specific aspects of the audience or the context, with the individuals often unaware of how deeply they are influenced by the social environment and their interpersonal objectives (Jones, 1990; Schlenker, 1980, 1985; Schlenker and Wowra, 2003; Tetlock and Manstead, 1985). Automatic self-presentation is most common in familiar environments among close acquaintances or friends, where individuals feel a sense of security and positive regard (e.g. while relaxing at home with closer friends) (Tyler, 2012). Self-presentation is also more likely to be automatic when it aligns with one’s established self-image and personality traits. Conversely, self-presentation often involves controlled processes in situations where the stakes or the audience are significant or where the individual is unsure about the impression they wish to project (Schlenker et al., 1996; Schlenker and Pontari, 2000). Under such conditions, individuals tend to deliberately focus on the impression they are making, meticulously planning and rehearsing their behavior (Tice et al., 1995; Schlenker et al., 1996). When required to deviate from their automatic mode of self-presentation, individuals may need to exert more cognitive effort, which can detract from their performance or quality of life (Vohs et al., 2005; Schlenker, 2003).

### v2

@@ Abstract.
Adolescents are particularly vulnerable to the pressures created by social media, such as heightened self-consciousness and the need for extensive self-presentation. In this study, we investigate how BeReal, a social media platform designed to counter some of these pressures, influences adolescents’ self-presentation behaviors. We interviewed 29 users aged 13-18 to understand their experiences with BeReal. We found that BeReal’s design focuses on spontaneous sharing, including randomly timed daily notifications and reciprocal posting, discourages staged posts, encourages careful curation of the audience, and reduces pressure on self-presentation. The space created by BeReal offers benefits such as validating an unfiltered life and reframing social comparison, but its approach to self-presentation is sometimes perceived as limited or unappealing and, at times, even toxic. Drawing on this empirical data, we propose design guidelines for platforms that support authentic self-presentation while fostering reciprocity and expanding beyond spontaneous photo-sharing. These guidelines aim to enable users to portray themselves more comprehensively and accurately, ultimately supporting teens’ developmental needs, particularly in building authentic relationships.
@@ 1. Introduction
The prevalence of social isolation has increased dramatically in recent decades. The number of people reporting no confidants for discussing important matters has tripled, while the average person’s social network has shrunk by one-third (McPherson et al., 2006). By 2004, nearly 25% of Americans reported having no confidant, up from 10% in 1985 (McPherson et al., 2006). This decline in social connections is concerning, as relationships significantly impact physical and mental health, longevity, and happiness (Cacioppo and Cacioppo, 2014; House et al., 1988; Cohen, 2004). The 75-year Study of Adult Development found that strong relationships are key to happiness and health (Waldinger and Schulz, 2016). A 2010 meta-analysis of over 300,000 participants revealed that weak social ties are as harmful to health as alcoholism (Holt-Lunstad et al., 2010). For adolescents, the need for trusted interpersonal connections, especially with peers, is even more crucial than for adults (Erikson, 1994a; Davis, 2012). With 95% of teens aged 13-17 actively engaging with friends on social media (noa, 2023), the nature of these social interactions has fundamentally changed. As peer relationships increasingly occur online, it’s essential to evaluate how technology supports teens’ relationship-building.
@@
Social Penetration Theory (Altman and Taylor, 1973) explains that relationships develop through reciprocal and authentic self-disclosure, deepening over time (Carpenter and Greene, 2015). However, social media often provides what has been termed “empty calories of social interaction” (Prinstein, 2023). While these interactions may seem to meet adolescents’ developmental needs, they often lack the substance necessary for genuine social benefits. Social media use can in fact exacerbate adolescents’ developmental vulnerabilities, such as heightened sensitivity to social feedback and increased emotional reactivity due to ongoing brain maturation (Somerville, 2013a; Steinberg, 2005; Falk et al., 2014; Peake et al., 2013; Somerville, 2013b). This can hinder the formation of meaningful, supportive relationships. Despite these shortcomings, most teens regularly engage in extensive self-presentation on social media, sharing carefully curated content, an experience that can induce anxiety and distort users’ perceptions of reality (Aacap, [n. d.]).
@@
BeReal is a social media platform designed to address the challenge of fostering genuine relationships online, where authentic interactions are often scarce. BeReal has a specific focus: “everyday get a notification to share a genuine glimpse into your real life with the people you care most about” which they believe enables users to “show your friends who you really are” (noa, [n. d.]). Though BeReal was launched in 2019, long before the U.S. Surgeon General’s 2023 advisory on social media and youth mental health (noa, 2023), its design and mission align with the advisory’s recommendation that tech companies “seek to maximize the potential benefits and avoid design features that attempt to maximize time, attention, and engagement” (noa, 2023).
@@
To answer these questions, we conducted semi-structured interviews with 29 BeReal users aged 13 to 18, asking about their general experiences with BeReal. We found that BeReal’s features and affordances are effective in encouraging users to share casual, uncurated snapshots of their daily lives. Features such as randomly timed daily notifications encourage authentic self-presentation via unstaged, unfiltered photo-sharing, and design elements like reciprocal posting encourage users to curate their audience, thereby allowing their self-presentation process to be less intentional or controlled. The participants saw many benefits of such designs. They especially appreciated that BeReal nudged users to collectively share more casual and less curated content, a change they said they sought on social media. However, some participants also felt that their ability to present themselves accurately and intentionally was too limited, and they explained that unfiltered photos taken at unplanned times are just one way of giving others insight into their lives. Others said that BeReal’s explicit prioritization of “authenticity” in self-presentation sometimes led to toxic competition and judgment about which users are truly being authentic.
In this paper, we provide empirical evidence of BeReal’s success in encouraging adolescent users to present themselves authentically. This serves as a proof of concept that design can significantly influence self-presentation. We also document the limitations of BeReal’s approach and outline design recommendations for creating a platform that encourages developmentally supportive online self-presentation for adolescents, such as scaffolding reciprocal sharing. Based on our findings, we introduce the Social Media Self-Presentation Matrix (SMSM). Grounded in theories on self-presentation goals and processes (Schlenker, 2003), this matrix presents ways that social media design can support different self-presentation strategies. It does so through four facets of self-presentation, spanning axes from Automatic to Controlled and from Self-Enhancing to Self-Expressive. Using the matrix, we highlight an untapped design opportunity to support Controlled and Self-Expressive self-presentation, which could offer an alternative way for social media platforms to achieve the goals initially intended by BeReal. Further, we propose a design agenda to support better relationship-building online based on authenticity for teens on social media. We hope this work will encourage designers and researchers to more comprehensively support authentic exchanges in teen peer-relationship-building on social media, addressing a developmental need that is currently unmet.
@@ 2.1. Adolescents and the Developmental Significance of Authentic Self-Disclosure
Forming close, trusting relationships is fundamental to people’s sense of meaning, happiness, and long-term well-being (Argyle, 2013). For adolescents, this intrinsic need for trusted interpersonal connections, particularly with peers, is even more pronounced than it is for adults (Erikson, 1994b; Davis, 2012). As described by the Social Penetration Theory (SPT) (Altman and Taylor, 1973), interpersonal relationships develop from initial, surface-level interactions to deeper, more intimate connections through a gradual process of self-disclosure. The breadth and depth of these disclosures increase as the relationship strengthens and matures. Also central to this progression is the norm of reciprocity, which posits that mutual exchange is fundamental to these evolving relationships.
@@
For self-disclosure to effectively enhance intimacy and trust in relationships, it must be authentic. Authenticity—the characteristic of being aligned with one’s intrinsic cues rather than being influenced by external expectations or pressures (Peets and Hodges, 2018)—ensures that the information shared truly represents the individual’s self, fostering genuine connections and understanding between people (Laurenceau et al., 1998; Peets and Hodges, 2018). Expressing their authentic selves and receiving acknowledgment and affirmation from peers is crucial for individuals, particularly adolescents, to develop a clear and coherent self-concept (Gardner, 2013). The formation of a coherent identity, rooted in an authentic self that is recognized and validated by one’s social circle, is fundamental to adolescent development (Erikson, 1994a). Research has shown that authenticity in self-disclosure is associated with life satisfaction (Bailey et al., 2020), particularly among girls, linking authenticity to self-esteem and overall well-being (Impett et al., 2008). Furthermore, a higher degree of authenticity in self-disclosing communication is connected to enhanced self-perceptions, reduced feelings of isolation in school, and greater satisfaction in relationships among adolescents (Peets and Hodges, 2018).
@@
In recent years, particularly exacerbated by the COVID-19 pandemic and subsequent physical distancing measures, adolescents have increasingly turned to social media for social connection (Hamilton et al., 2022; Pitt et al., 2021). As of 2022, nearly all teenagers aged 13-17 (95%) own a smartphone or access the internet (97%), with nearly half (46%) using the internet “almost constantly” (Vogels et al., 2022). Previous studies have identified that the primary motivation behind teens’ use of digital media is to maintain connections with existing friends (Ito, 2013; Boyd, 2008a; Davis, 2012). Given the critical role of online platforms in adolescent lives and the value placed on authenticity in their relationships, it is understandable that modern teenagers actively seek exchanges of authentic self-disclosure on social media and broader media contexts (Darr and Doss, 2022; Stephanie Rivas-Lara Becky Pham Jasmine Baten Annie Meyers Yalda T. Uhls, 2022). Further, the significantly reduced social cues on online platforms heighten uncertainty, making self-disclosure especially crucial for (adolescent) users to understand each other and strengthen their connections (Schlosser, 2020; Tidwell and Walther, 2002). This underscores the importance of examining how technology can support or undermine the development and maintenance of these relationships, with a particular focus on the reciprocal exchange of authentic self-disclosure throughout the stages of social penetration.
@@ 2.2. Authentic Self-Disclosure and Self-Presentation
Given that self-disclosure, especially in online contexts, involves extensive interpersonal interactions, it can be likened to a performance (Goffman, 2016) or an exhibition (Hogan, 2010) that symbolically communicate information about oneself to others (Blumer, 1986). As a result, self-presentation becomes an inherent part of this process. Self-disclosure is defined as “verbally communicating personal information about the self to another person” (Forgas, 2011) (p. 449), and reflects communicating a factual representation of oneself, regardless of its effect on one’s public image. In contrast, self-presentation is defined as “the goal-directed activity of controlling information to influence the impressions formed by an audience about the self” (Schlenker and Wowra, 2003) (p. 871). Thus, self-disclosure typically aligns with authenticity, whereas self-presentation may suggest a lack of authenticity. However, these two are not mutually exclusive (Johnson, 1981); revealing true aspects of oneself can still be compatible with self-presentation, as such revelations can also support a favorable public persona (Schlosser, 2020). We adopt the view in this paper that self-presentation is a strategic form of self-disclosure that predominates on social media, where users, acting as performers, anticipate positive outcomes from their disclosures.
@@
Although self-presentation is often perceived as strategic gamesmanship, this view does not capture its full complexity (Schlenker, 2003). In reality, self-presentation can be deeply “authentic,” as in, it typically involves a portrayal that is somewhat enhanced and idealized, yet genuinely believed by the person presenting it (Brown, 2014; Greenwald and Breckler, 1985; Williams and Gilovich, 2008). Self-presentation often aims for accurate self-portrayal (Baumeister, 1982; Cheek and Hogan, 1983; Leary, 2019; Schlenker, 1980; Schlenker and Pontari, 2000). This approach may serve several purposes: to ensure others perceive one in a manner that elicits respect and treatment congruent with one’s identity; to receive affirming feedback that addresses personal uncertainties; and to embody the principle that “honesty is the best policy,” fostering a sense of authenticity while mitigating risks associated with deception. Furthermore, intentional self-presentation does not inherently signify inauthenticity; Schlenker (Schlenker, 2003) uses the analogy of how a skilled lecturer meticulously crafts their presentation to effectively convey genuine and significant information to their audience. This stands in stark contrast to a disorganized lecturer whose lack of preparation may come across as insincere or self-absorbed. Similarly, the careful arrangement of personal information may better reflect an authentic self-depiction as perceived by the individual) (Schlenker and Pontari, 2000). In other words, presenting accurate information requires as much skill in self-presentation as presenting inaccurate information if it is to achieve the intended effect on the audience.
@@
Authentic self-presentation can manifest through either automatic or controlled cognitive processes. Automatic processes are those which 1) occur without conscious awareness, 2) require minimal cognitive effort, 3) do not require conscious monitoring, and 4) are initiated involuntarily without outside triggers (Schlenker, 2003; Bargh, 1996). In such instances, self-presentation scripts may be triggered automatically by specific aspects of the audience or the context, with the individual often unaware of how deeply they are influenced by the social environment and their interpersonal objectives (Jones, 1990; Schlenker, 1980, 1985; Schlenker and Wowra, 2003; Tetlock and Manstead, 1985). Automatic self-presentation is most common in familiar environments among close acquaintances or friends, where individuals feel a sense of security and positive regard (e.g., while relaxing at home with closer friends) (Tyler, 2012; Schlenker, 2003). Self-presentation is also more likely to be automatic when it aligns with one’s established self-image and personality traits (Schlenker, 2003). Conversely, self-presentation often involves controlled processes in situations where the stakes or the audience are significant or where the individual is unsure about the impression they wish to project (Schlenker et al., 1996; Schlenker and Pontari, 2000). Under such conditions, individuals tend to deliberately focus on the impression they are making, meticulously planning and rehearsing their behavior (Tice et al., 1995; Schlenker et al., 1996). The caveat of having to deviate from their automatic mode of self-presentation is that individuals may need to exert more cognitive effort, which can detract from their performance or quality of life (Vohs et al., 2005; Schlenker, 2003).